OpenStudy (anonymous):

rewrite equation in slope intercept y=-x+6 x+y=20

OpenStudy (anonymous):

Is this a system or are they separately being put into slope-intercept form?

OpenStudy (anonymous):

a system i guess

OpenStudy (anonymous):

Ok so if it's a system then are you given a certain way to solve like elimination method or substitution method because I can do both and get your answer for you.

OpenStudy (anonymous):

it also says determine whether the lines are parallel

OpenStudy (anonymous):

Ok, if you're just rewriting the equations in slope-intercept then the asnwer is y=-x+6 (because it's already in slope-intercept form) y= -x + 20 (Because you wanted to get why by itself by subtracting X on both sides) So, since both of the slopes to this are technically -1 (since there is an "invisible " 1 in front of the x) these lines are parallel If you're solving the system you would use substitutio method because it's easiest So x(-x+6) =20 because I plugged the first equation into the Y of the second equation since y=-x+6 The math it out x(-x+6)=20 -2x+6x=20 4x+20 x=5 Plug X back into either equation y=-(5)+6 y=1 So the answer to solving the equation is (5,1)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome